what is the slope of (-2,-8), (1,4)
Answer by haileytucki(390)

(-2,-8),(1,4)
Slope is equal to the change in y over the change in x, or 'rise over run'.
m=(change in y)/(change in x)
The change in x is equal to the difference in x-coordinates (also called run), and the change in y is equal to the difference in y-coordinates (also called rise).
m=(y2-y1)/(x2-x1)
Substitute in the values of x and y into the equation to find the slope.
m=(4-(-8))/(1-(-2))
Multiply -1 by each term inside the parentheses.
m=(4-(-8))/(1+2)
Add 2 to 1 to get 3.
m=(4-(-8))/(3)
Multiply -1 by each term inside the parentheses.
m=(4+8)/(3)
Add 8 to 4 to get 12.
m=(12)/(3)
Reduce the expression (12)/(3) by removing a factor of 3 from the numerator and denominator.
m=4